# On the electronic normal modes of the Meyer–Miller–Stock–Thoss representation of non-adiabatic dynamics

### v1

@@ III.3.1 In mapping variables
The CF in Eqn. (19) can be obtained in terms of mapping variables, where the real and imaginary terms are outlined in the Section IA of the Supplementary Material and follows a similar derivation as in references [ 63] and [ 53], but neglecting the nuclear terms. This results in the real-time term being,
@@
where which is the symmetrized operator. We use the shorthand notation to describe the following as the operators only act once in the path-integral loop,
@@
where . The CF can be calculated by averaging over many trajectories using the sampling outlined in Section IB of the Supplementary Material. If any imaginary terms arise during computation, these can be neglected as the correlation function is proven to be real in Section IC of the Supplementary Material.
@@ III.4 Conservation of the quantum Boltzmann distribution
In Appendix D, we prove QBD conservation for both a single-bead and multi-bead calculation for a single trajectory. In the multi-bead case, acting the Liouvillian brings down a either side that bead’s Boltzmann term, as seen in Fig. D.1. This means that the terms only cancel out when completing the path-integral loop, so all beads are required for the QBD to be conserved. The multi-bead proof can be used to prove that a calculation with all the normal modes included will also conserve the distribution. A back-transformation can be performed to obtain the bead-form from the normal-mode calculation, and this is equivalent to including all the beads. This also suggests that truncating in beads or normal modes will not conserve the QBD for a single trajectory, likely a consequence of the observable not being a function of only a finite number of the lowest normal modes. We investigate this computationally later in this article and leave further algebraic investigation of this as future work.
@@ IV Results
For simplicity, we consider only the electronic system in our calculation, as mentioned earlier. We think it is very unlikely that a general QBD-conserving method will be found which does not conserve the QBD for the electronic-only system. Here, we discuss the first electronic state population auto-CF and the conservation of the QBD when truncating in beads and normal modes. We also provide additional plots that aid our understanding of how each bead and normal mode contributes to the calculations.
@@
where the asymmetry, and the electronic coupling, , such that we model an asymmetric system. Results for a symmetric system where can be seen in the Supplementary Material. All the results shown here are for a 8 bead calculation where we have also transformed into the 8 normal modes and truncated to the lowest 1, 3 and 5 beads and normal modes.
@@ IV.1 Autocorrelation function
Firstly, we calculate the CF using all 8 beads and normal modes of an 8-bead calculation, and after truncating in beads or normal modes, shown in Figure 1. When truncating, we keep the beads/normal modes that are truncated at the zero-time sampled value. We see that the more beads/normal modes that are included, the result systematically becomes more accurate compared to the exact KT result (solid black line), as seen in Figure 1. This plot demonstrates numerically what we have shown algebraically in Eqn. (40) that the observable is a function of not only the lowest normal modes but all normal modes, and that the higher normal modes do still contribute significantly to the dynamics.
@@ IV.2 Quantum Boltzmann Distribution Conservation
The single trajectory results in Figure 2 agree with the algebra in Appendix D. The full 8 bead and normal mode calculations conserve the QBD, as seen by the flat line (green dotted circle and dashed), for a single trajectory. The dotted circle and dashed green lines are the same such that it is only possible to see the dotted circle green line in this plot. Any truncation in both the beads and normal modes results in the line oscillating, indicating that this term is not conserved. The energy of a single trajectory, , can also be tested for conservation (Figure S.2 in the Supplementary Material) where we see the same trend as for the Boltzmann term.
We can test the conservation of an ensemble of trajectories by calculating the expectation value through the evolved CF, . Although algebraically we can show that all beads are required to obtain QBD conservation, this is challenging to determine in normal modes as the transformation matrix cannot be eliminated. While we can show that including all normal modes conserves the QBD, we leave a full description of the effect of truncation algebraically as further work. The numerical results for a single trajectory show that truncating in normal modes does not conserve the QBD. However, this does not mean there will not be an averaging effect that results in conservation for many trajectories.
@@ IV.3 MMST normal mode distribution
To try and understand why all the normal modes need to be included to obtain an accurate CF, we look at the distribution of MMST normal modes. In Figure 5, we compare the distribution of MMST normal modes of (for the first electronic state using the asymmetric potential outlined earlier) with nuclear position normal modes for a ring-polymer in a harmonic potential, . For the ring-polymer normal modes (Figure 5a), we see the distribution significantly narrows and becomes taller as the mode index increases. This is as the normal modes are constrained by the springs between the beads. In Matsubara dynamics, the higher, truncated (unevolved) normal modes can be integrated out by performing a contour integral which is valid at .Hele et al. (2015b, a) This results in spring-like terms between the higher normal modes and so the truncated normal modes will be constrained in a similar manner to the ring-polymer normal modes seen here.Hele et al. (2015b, a) However, for the MMST normal modes (Figure 5b), the distributions are all the same and we do not see any constraint.
In the Supplementary Material, we investigate several other metrics involving MMST variables, such as the total and state populations, as seen in Supplementary Material Figures S.5-7. However, none of these metrics resulted in a narrowing of the normal mode distributions. This shows that normal modes of MMST variables, and of the associated population metrics based on MMST variables, are not a good metrics as there is no constraint of higher normal modes and we do not obtain accurate dynamics or trajectory-level QBD conservation after truncation.
@@ IV.4 Summary
Criterion | Nuclear Matsubara Modes | MMST Electronic Modes |
Observable function of finite number of NM | ✓ | ✗ |
Constrained higher NM | ✓ | ✗ |
Single trajectory QBD conservation upon truncation | ✓ | ✗ |
Ensemble QBD conservation upon truncation | ✓ | ✓(numerically) |
@@
Although it appears that the MMST representation is not optimal for describing electronic normal modes, this does not mean that no other metric exists which might be suitable. Truncating in electronic normal modes described by some other metric may still allow us to obtain both accurate CFs from a finite number of normal modes and conservation of the QBD. As mentioned earlier, we intend to investigate other metrics, such as spin-mapping normal modes, for these properties in future work.
@@ V Conclusions
In this article we have considered in detail for the first time the electronic normal modes of the position and momentum variables of the Meyer–Miller–Stock–Thoss (MMST) representation of non-adiabatic dynamics. We have discussed the use of electronic normal modes calculated through the ‘obvious’ method of directly transforming and from the MMST mapping Hamiltonian, as is done for the nuclear variables in Matsubara dynamics. Whilst we could have considered an alternative transformation matrix or a different metric to take normal modes of, we believe the study of MMST normal modes with the same transformation as for ring-polymer normal modes of significant interest. The transformation is well understood and utilised in Matsubara dynamics,Ceriotti et al. (2010); Hele et al. (2015a) and the proposed NA-Mats method is constructed in MMST variables but does not take electronic normal modes or generally conserve the QBD.Chowdhury and Huo (2021)
@@
We do not find that the observables are a function of only a finite number of the lowest normal modes, that the higher normal modes are constrained, that the QBD is conserved for a single trajectory when truncating in either beads or normal modes, or that accurate dynamics are obtained from truncation. The ensemble calculations are also particularly challenging to converge with a 2-level electronic-only system and there is an averaging effect that results in QBD conservation for all tested truncations. These results suggest that the MMST variables may not be the best metric to derive a non-adiabatic dynamics method satisfying the four criteria outlined in the introduction. However, this does not mean that we cannot find a non-adiabatic Matsubara method that will conserve the QBD distribution, for both a single and ensemble of trajectories, but rather that it requires a different metric from the MMST and . This also suggests that were one to construct a general non-adiabatic Matsubara method that truncated in both the electronic and nuclear normal modes it would be unlikely to conserve the QBD (at least for a single trajectory) and its dynamics would be unlikely to be a good approximation to the exact quantum result. These results may also explain why, despite the variety of non-adiabatic dynamics methods using the MMST representation, we are unaware of any such method which conserves the QBD.Meyer and Miller (1979); Stock and Thoss (2005); Church et al. (2018); Richardson et al. (2017); Sun, Wang, and Miller (1998a, b); Ananth, Venkataraman, and Miller (2007); Kelly et al. (2012); Saller, Kelly, and Richardson (2020); Huo and Coker (2011); Kirrander and Vacher (2020)
In future research we intend to explore alternative metrics for electronic normal modes, including spin-mapping,Runeson and Richardson (2019, 2020); Bossion, Chowdhury, and Huo (2021); Bossion et al. (2022) as this may lead to a simpler correlation function that is easier to converge and for which observables can be written as a finite function of only the lowest normal modes, and the higher normal modes are constrained such that truncation results in an accurate correlation function and QBD conservation.
@@ Acknowledgements
The data that support the findings of this study are openly available in UCL Research Data Repository at http://doi.org/[to be inserted], reference number [to be inserted].
@@ D.1 Single bead
For 1 bead, the Liouvillian is Eqn. (18) (where we let ) and the CF is Eqn. (36) in the single bead limit such that,
@@
where we have used the fact that and commute to reverse the order of these matrices. The two terms we obtain per bead can be seen as the Liouvillian operator ‘pulling’ down a to the right of the Boltzmann term for that bead and a to the left of the Boltzmann term for that bead, Figure D.1.

### v2

@@ III.3.1 In mapping variables
The CF in Eqn. (19) can be obtained in terms of mapping variables, where the real and imaginary terms are outlined in the Section IA of the Supplementary Material and follows a similar derivation as in references [ 66] and [ 56], but neglecting the nuclear terms. This results in the real-time term as,
@@
where which is the symmetrized operator. The above is the short-hand notation to describe the following as the operators only act once in the path integral loop,
@@
where . The CF can be calculated by averaging over many trajectories using the Metropolis Monte-Carlo sampling outlined in Section IB of the Supplementary Material. If any imaginary terms arise during computation, these can be neglected as the correlation function is proven to be real in Section IC of the Supplementary Material.
@@ III.4 Conservation of the quantum Boltzmann distribution
In Appendix D, we prove QBD conservation for both a single-bead and multi-bead calculation for a single trajectory. In the multi-bead case, acting the Liouvillian brings down a either side that bead’s term, as seen in Figure 7. This means that the terms only cancel out when completing the path integral loop, so all beads are required for the QBD to be conserved. The multi-bead proof can be used to prove that a calculation with all the normal modes included will also conserve the distribution. A back-transformation can be performed to obtain the bead form from the normal mode calculation, and this is equivalent to including all the beads. This also suggests that truncating in beads or normal modes will not conserve the QBD for a single trajectory, likely a consequence of the observable not being a function of only a finite number of the lowest normal modes and that the transformation matrix into normal modes cannot be removed when truncating. We investigate this computationally later in this article and leave further algebraic investigation of this as future work.
@@ IV Results
For simplicity, we consider only the electronic system in our calculation, as mentioned earlier. We think it is very unlikely that a general QBD-conserving method will be found which does not conserve the QBD for the electronic-only system. Here, we discuss the first electronic state population auto-CF and the conservation of the QBD when truncating in beads and normal modes. We also provide additional plots that aid our understanding of how each bead and normal mode contributes to the calculations. When truncating in the bead representation ( and ), we propagate the first beads and keep the remaining beads at their initial time values. Likewise, when truncating in normal modes ( and ), we propagate the lowest normal modes and all higher normal modes are kept at their initial time values. We can do this as the normal modes are separable for the electronic-only system, seen in Eqn. (29). A schematic for this can be seen in Figure 1.
@@
where the asymmetry, and the electronic coupling, , such that we model an asymmetric system. In all cases, we utilised reduced units such that . All the results shown here are for a 8 bead calculation where we have also transformed into the 8 normal modes and truncated to the lowest 1, 3 and 5 beads and normal modes. Results for a symmetric system where with and can be seen in the Supplementary Material (Figures S.1–6) and for an asymmetric potential where (Figures S.7–9). We note that an 8 bead calculation considering 5 normal modes is significantly outside the Matsubara limit (where the number of beads is far greater than the number of normal modes), however, larger numbers of beads makes convergence challenging due to the Metropolis Monte-Carlo sampling and we find provides us with enough information to probe adequately the properties of the MMST normal modes.
@@ IV.1 Autocorrelation function
Firstly, we calculate the CF using a full 8 bead and normal mode calculation, and when truncating in beads or normal modes, shown in Figure 2. When truncating, we keep the truncated beads or normal modes at the zero-time sampled value. We see that the more beads/normal modes that are included, the result systematically becomes more accurate compared to the exact KT result (solid black line), as seen in Figure 2. This plot demonstrates numerically what we have shown algebraically in Eqn. (40) that the observable is a function of not only the lowest normal modes but all normal modes, and that the higher normal modes do still contribute significantly to the dynamics. Each set of bead and normal mode results converge on similar values, highlighted by the results in the Supplementary Material (Figure S.2) where convergence is easier to obtain, emphasizing that the transformation does not change the dynamics for the electronic-only system.
@@ IV.2 Quantum Boltzmann Distribution Conservation
The single trajectory results in Figure 3 agree with the algebra in Appendix D. The full 8 bead and normal mode calculations conserve the QBD, as seen by the flat line (green dotted and dashed), for a single trajectory. The dotted and dashed green lines are the same such that it is only possible to see the dashed green line in this plot. Any truncation in both the beads and normal modes results in the line oscillating, indicating that this term is not conserved. The energy of a single trajectory, , can also be tested for conservation (Figure S.3 in the Supplementary Material) where we see the same trend as for the Boltzmann term.
We can test the conservation of an ensemble of trajectories by calculating the expectation value through the evolved CF, . Although algebraically we can show that all beads are required to obtain QBD conservation for a single trajectory, this is challenging to determine in normal modes due to the inability to transform between in beads and in normal modes when truncating. This makes it unlikely that the QBD will be conserved when truncating in normal modes. While we can show that including all normal modes conserves the QBD, we leave a full description of the effect of truncation algebraically as further work. The numerical results for a single trajectory show that truncating in normal modes does not conserve the QBD. However, this does not mean there will not be an averaging effect that results in conservation for many trajectories.
@@ IV.3 MMST normal mode distribution
To try and understand why all the normal modes need to be included to obtain an accurate CF, we look at the distribution of MMST normal modes. In Figure 6, we compare the distribution of MMST normal modes of (for the first electronic state using the asymmetric potential outlined earlier) with nuclear position normal modes for a ring-polymer in a harmonic potential, . For the ring-polymer normal modes (Figure 6a), we see the distribution significantly narrows and becomes taller as the mode index increases. This is as the normal modes are constrained by the springs between the beads. In Matsubara dynamics, the higher, truncated (unevolved) normal modes can be integrated out by performing a contour integral which is valid at .Hele et al. (2015b, a) This results in spring-like terms between the higher normal modes and so the truncated normal modes will be constrained in a similar manner to the ring-polymer normal modes seen here.Hele et al. (2015b, a) However, for the MMST normal modes (Figure 6b), the distributions are all the same, due to a lack of springs between the electronic DoF, and we do not see any narrowing.
In the Supplementary Material, we investigate several other electronic population metrics involving MMST variables as seen in Figures S.11-13. However, none of these metrics resulted in a narrowing of the normal mode distributions. This shows that normal modes of MMST variables, and of the associated population metrics based on MMST variables, are not optimal metrics for constructing a QBD-conserving method with a Matsubara approach.
@@ IV.4 Summary
| Criterion |
| Observable function of finite number of NM | ✓ | ✗ | ||||
| Constrained higher NM | ✓ | ✗ | ||||
| Truncated single trajectory QBD conservation | ✓ | ✗ | ||||
✓ | ✓(numerically) | ||||
@@
Although it appears that the MMST representation is not optimal for describing electronic normal modes, due to the lack of springs between electronic DoF, this does not mean that no other metric exists which might be suitable. Truncating in electronic normal modes described by some other metric may still allow us to obtain both accurate CFs from a finite number of normal modes and conservation of the QBD. We leave investigation of alternative metrics as future research.
@@ V Conclusions
In this article, we have considered in detail for the first time the electronic path integral normal modes of the position and momentum variables of the Meyer–Miller–Stock–Thoss (MMST) representation of nonadiabatic dynamics. We have discussed the use of electronic normal modes calculated through the ‘obvious’ method of directly transforming and from the MMST mapping Hamiltonian, as is done for the nuclear variables in Matsubara dynamics. Whilst we could have considered an alternative transformation matrix or a different metric to take normal modes of, we believe the study of MMST normal modes with the same transformation as for ring-polymer normal modes is of significant interest. The transformation is well understood and utilised in Matsubara dynamics,Ceriotti et al. (2010); Hele et al. (2015a) and the proposed NA-Mats method is constructed in MMST variables but does not take electronic normal modes or generally conserve the QBD.Chowdhury and Huo (2021)
@@
We do not find that the observables are a function of only a finite number of the lowest normal modes, that the distribution of the higher normal modes significantly narrows, that the QBD is conserved for a single trajectory when truncating in either beads or normal modes, or that accurate dynamics are obtained from truncation. The ensemble calculations are also particularly challenging to converge with a 2-level electronic-only system and there is an averaging effect that results in QBD conservation for all tested truncations. The lack of narrowing of the distribution arises from the lack of a Boltzmann constraint or penalty in these electronic DoF, unlike in the conventional nuclear ring-polymer where the imaginary-time path integral leads to ‘springs’ which results in a narrowing in the distribution of higher normal modes. These results suggest that normal modes of the MMST variables may not be the best metric to derive a nonadiabatic dynamics method satisfying the four criteria outlined in the introduction. However, this does not mean that we cannot find a nonadiabatic Matsubara method that will conserve the QBD distribution, for both a single and ensemble of trajectories, but rather that it requires a different metric from the MMST and . Whilst here we consider a few simple metrics constructed from MMST electronic variables, this is by no means a complete search and we intend to consider additional metrics as future research, including spin-mapping,Runeson and Richardson (2019, 2020); Bossion, Chowdhury, and Huo (2021); Bossion et al. (2022) which may result in a QBD conserving method.
This also suggests that were one to construct a general nonadiabatic Matsubara method based on MMST variables that truncated in both the electronic and nuclear normal modes it would be unlikely to conserve the QBD (at least for a single trajectory) and its dynamics would be unlikely to be a good approximation to the exact quantum result. These results may also explain why, despite the variety of nonadiabatic dynamics methods using the MMST representation, we are unaware of any such method which uses the MMST Hamiltonian to propagate dynamics and also conserves the QBD.Meyer and Miller (1979); Stock and Thoss (2005); Church et al. (2018); Richardson et al. (2017); Sun, Wang, and Miller (1998a, b); Ananth, Venkataraman, and Miller (2007); Kelly et al. (2012); Saller, Kelly, and Richardson (2020); Huo and Coker (2011); Kirrander and Vacher (2020)
Supplementary Material
The supplementary material contains additional algebra for: the CF derivation in mapping variables, the partition function and sampling details and, proof that the CF is real. We provide additional figures for a symmetric potential with and beads. For the asymmetric potential we present: the results, ensemble convergence results with different truncations and, the distributions of alternative electronic population metrics (based on MMST variables).
@@ Acknowledgements
The data that support the findings of this study are openly available in UCL Research Data Repository at https://doi.org/10.5522/04/c.7993391, reference number 7993391.
@@ D.1 Single bead
For 1 bead, the Liouvillian is Eqn. (18) (where we let ) and the CF is Eqn. (III.3.1) in the single bead limit such that,
@@
where we have used the fact that and commute to reverse the order of these matrices. The two terms we obtain per bead can be seen as the Liouvillian operator ‘pulling’ down a to the right of the Boltzmann term for that bead and a to the left of the Boltzmann term for that bead.